\begin{thebibliography}{10}

\bibitem{Akhmerov-Kurbatov}
R.~R. Akhmerov and V.~G. Kurbatov, \emph{Exponential dichotomy and stability of
  neutral type equations}, J. Differential Equations \textbf{76} (1988), no.~1,
  1--25. \MR{964610}

\bibitem{BaskakovMS15:eng}
A.~G. Baskakov, \emph{Estimates for the {G}reen's function and parameters of
  exponential dichotomy of a hyperbolic operator semigroup and linear
  relations}, Mat. Sb. \textbf{206} (2015), no.~8, 23--62, (in Russian);
  English translation in \emph{Sb. Math.}, \textbf{206} (2015), no.~8,
  1049--1086. \MR{3438589}

\bibitem{Boichuk-Pokutnii}
A.~A. Boichuk and A.~A. Pokutnii, \emph{Bounded solutions of linear
  differential equations in a {B}anach space}, Nonlinear Oscillations
  \textbf{9} (2006), no.~1, 3--14. \MR{2369770}

\bibitem{Bulgakov80:eng}
A.~Ya. Bulgakov, \emph{An effectively calculable parameter of stability quality
  of systems of linear differential equations with constant coefficients},
  Sibirsk. Mat. Zh. \textbf{21} (1980), no.~3, 32--41, 235, (in Russian);
  English translation in \emph{Siberian Mathematical Journal},
  \textbf{21}(1980), no. 3, 339--347. \MR{574960}

\bibitem{Bulgakov89a:eng}
\bysame, \emph{An estimate of the {G}reen matrix and the continuity of the
  dichotomy parameter}, Sibirsk. Mat. Zh. \textbf{30} (1989), no.~1, 178--182,
  (in Russian); English translation in \emph{Siberian Mathematical Journal},
  \textbf{30}(1989), no. 1, 139--142. \MR{995033}

\bibitem{Chicone-Latushkin99}
C.~Chicone and Y.~Latushkin, \emph{Evolution semigroups in dynamical systems
  and differential equations}, Mathematical Surveys and Monographs, vol.~70,
  American Mathematical Society, Providence, RI, 1999. \MR{1707332}

\bibitem{Daletskii-Krein:eng}
Ju.~L. Dalecki{\u\i} and M.~G. Kre{\u\i}n, \emph{Stability of solutions of
  differential equations in {B}anach space}, Translations of Mathematical
  Monographs, vol.~43, American Mathematical Society, Providence, RI, 1974.
  \MR{0352639}

\bibitem{Gelfond:eng}
A.~O. Gel$'$fond, \emph{Calculus of finite differences}, International
  Monographs on Advanced Mathematics and Physics, Hindustan Publishing Corp.,
  Delhi, 1971, Translation of the third Russian edition. \MR{0342890}

\bibitem{Gil-LMA-93}
M.~I. Gil', \emph{Estimate for the norm of matrix-valued functions}, Linear and
  Multilinear Algebra \textbf{35} (1993), no.~1, 65--73. \MR{1310964}

\bibitem{Godunov86:SMZh:eng}
S.~K. Godunov, \emph{The problem of the dichotomy of the spectrum of a matrix},
  Sibirsk. Mat. Zh. \textbf{27} (1986), no.~5, 24--37, 204. \MR{873707}

\bibitem{Godunov94:ODE:eng}
\bysame, \emph{Ordinary differential equations with constant coefficient},
  Translations of Mathematical Monographs, vol. 169, American Mathematical
  Society, Providence, RI, 1997, Translated from the 1994 Russian original.
  \MR{1465434}

\bibitem{Golub-Van_Loan96:eng}
G.~H. Golub and Ch.~F. Van~Loan, \emph{Matrix computations}, third ed., Johns
  Hopkins Studies in the Mathematical Sciences, Johns Hopkins University Press,
  Baltimore, MD, 1996. \MR{1417720}

\bibitem{Hartman02:eng}
Ph. Hartman, \emph{Ordinary differential equations}, second ed., Classics in
  Applied Mathematics, Society for Industrial and Applied Mathematics (SIAM),
  Philadelphia, PA, 2002. \MR{1929104}

\bibitem{Henry81:eng}
D.~Henry, \emph{Geometric theory of semilinear parabolic equations}, Lecture
  Notes in Mathematics, vol. 840, Springer-Verlag, Berlin--New York, 1981.
  \MR{610244}

\bibitem{Jordan}
Ch. Jordan, \emph{Calculus of finite differences}, third ed., Chelsea
  Publishing Co., New~York, 1965. \MR{0183987}

\bibitem{Kato:eng}
T.~Kato, \emph{Perturbation theory for linear operators}, Classics in
  Mathematics, Springer--Verlag, Berlin, 1995, Reprint of the 1980 edition.

\bibitem{Kurbatov-Kurbatova-CMAM18}
V.~G. Kurbatov and I.~V. Kurbatova, \emph{Computation of {G}reen's function of
  the bounded solutions problem}, Computational Methods in Applied Mathematics
  (2018), no.~4, 673--685.

\bibitem{Kurbatov-Kurbatova-QTDS18}
\bysame, \emph{The {G}elfand--{S}hilov type estimate for {G}reen's function of
  the bounded solutions problem}, Qualitative Theory of Dynamical Systems
  \textbf{17} (2018), no.~3, 619--629.

\bibitem{Kurbatov-Kurbatova-triangular1}
\bysame, \emph{Green's function of the problem of bounded solutions in the case
  of a block triangular coefficient}, April 2018, arXiv: 1804.01022.

\bibitem{Kurbatov-Kurbatova-Oreshina}
V.~G. Kurbatov, I.~V. Kurbatova, and M.~N. Oreshina, \emph{Analytic functional
  calculus for two operators}, April 2016, arXiv: 1604.07393.

\bibitem{Malyshev91:eng}
A.~N. Malyshev, \emph{Introduction to numerical linear algebra}, Nauka,
  Novosibirsk, 1991, (in Russian).

\bibitem{Massera-Schaffer:eng}
J.~L. Massera and J.~J. Sch\"affer, \emph{Linear differential equations and
  function spaces}, Pure and Applied Mathematics, Vol. 21, Academic Press, New
  York--London, 1966. \MR{0212324}

\bibitem{Pankov90:eng}
A.~A. Pankov, \emph{Bounded and almost periodic solutions of nonlinear operator
  differential equations}, Mathematics and its Applications (Soviet Series),
  vol.~55, Kluwer Academic Publishers, Dordrecht, 1990, Translated from the
  1985 Russian edition. \MR{1120781}

\bibitem{Pechkurov12:eng}
A.~V. Pechkurov, \emph{Bisectorial operator pencils and the problem of bounded
  solutions}, Izv. Vyssh. Uchebn. Zaved. Mat. (2012), no.~3, 31--41, (in
  Russian); English translation in \emph{Russian Math. (Iz. VUZ)}, \textbf{56}
  (2012), no.~3, 26--35. \MR{3076516}

\bibitem{Van_Loan77}
Ch.~F. Van~Loan, \emph{The sensitivity of the matrix exponential}, SIAM J.
  Numer. Anal. \textbf{14} (1977), no.~6, 971--981. \MR{0468137}

\bibitem{Watson95:eng}
G.~N. Watson, \emph{A treatise on the theory of {B}essel functions}, Cambridge
  Mathematical Library, Cambridge University Press, Cambridge, 1995, Reprint of
  the second (1944) edition. \MR{1349110}

\bibitem{Wolfram}
S.~Wolfram, \emph{The {M}athematica book}, fifth ed., Wolfram Media, New York,
  2003.

\end{thebibliography}

\documentclass[12pt, reqno]{amsart}
\usepackage{amsmath, amsthm, amscd, amsfonts, amssymb, graphicx, color}
\usepackage[mathscr]{eucal}
\usepackage[bookmarksnumbered, colorlinks, plainpages]{hyperref}
\hypersetup{colorlinks=true,linkcolor=red, anchorcolor=green, citecolor=cyan, urlcolor=red, filecolor=magenta, pdftoolbar=true}

\newtheorem{theorem}{Theorem}

\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}
\theoremstyle{remark}
\newtheorem{remark}{Remark}

\DeclareMathAlphabet\mathoo{U}{eur}{b}{n}

\DeclareMathOperator{\Real}{Re}
\DeclareMathOperator{\heaviside}{\eta}

\begin{document}
\setcounter{page}{1}

\title[An estimate of Green's function]{An estimate of Green's function\\ of the problem of bounded solutions\\ in the case of a triangular coefficient}

\author{V.~G. Kurbatov}
 \address{Department of Mathematical Physics,
Voronezh State University\\ 1, Universitetskaya Square, Voronezh 394018, Russia}
\email{\textcolor[rgb]{0.00,0.00,0.84}{kv51@inbox.ru}}

\author{I.V. Kurbatova}
 \address{Department of Software Development and Information Systems Administration,
Vo\-ro\-nezh State University\\ 1, Universitetskaya Square, Voronezh 394018, Russia}
\email{\textcolor[rgb]{0.00,0.00,0.84}{la\_soleil@bk.ru}}

\subjclass{Primary 65F60; Secondary 97N50}

\keywords{bounded solutions problem, Green's function, estimate, Schur decomposition}

\date{\today}

\begin{abstract}
An estimate of Green's function of the bounded solutions problem for the ordinary differential equation $x'(t)-Bx(t)=f(t)$ is proposed.
It is assumed that the matrix coefficient $B$ is triangular.
This estimate is a generalization of the estimate of  the matrix exponential proved by Ch.~F. Van~Loan.
\end{abstract}

\maketitle

\section*{Introduction}\label{s:Introduction}
Any square complex matrix $A$ can be represented in the triangular Schur form $A=Q^{-1}BQ$, where $B$ is triangular and $Q$ is unitary. Moreover, the Schur decomposition is calculated effectively and constitutes the preliminary stage of many algorithms connected with spectral theory. Thus, in numerical analysis, having a matrix $A$, one may assume that its triangular form is known. In its turn, any triangular matrix $B$ can be easily represented as the sum $B=D+N$ of a diagonal matrix $D$ and a strictly triangular matrix $N$. In~\cite{Van_Loan77} it is established the estimate
\begin{equation}\label{e:Loan est}
\lVert e^{At}\rVert=\lVert e^{Bt}\rVert\le e^{\alpha t}\sum_{k=0}^{n-1}\frac{\lVert Nt\rVert^k}{k!},\qquad t\ge0,\tag{$*$}
\end{equation}
where the matrix $A$ has the size $n\times n$ and $\alpha=\max\{\,\Real\lambda:\,\lambda\in\sigma(A)\,\}$. For other estimates of $e^{At}$, see, e.g.,~\cite{Bulgakov80:eng,Gil-LMA-93}, \cite[ch. 9, \S~1]{Kato:eng}. In this paper, we establish an estimate similar to~\eqref{e:Loan est} for Green's function of the bounded solutions problem (Theorem~\ref{t:fin}). Different estimates for Green's function are obtained in~\cite{Bulgakov89a:eng,Godunov86:SMZh:eng,Kurbatov-Kurbatova-QTDS18}.

We consider the differential equation
\begin{equation}\label{e:eq}
x'(t)-Ax(t)=f(t),\qquad t\in\mathbb R,\tag{$**$}
\end{equation}
where $A$ is a square complex matrix.
The \emph{bounded solutions problem} is the problem of finding a bounded solution~$x$ that corresponds to a bounded free term $f$. The bounded solutions problem is closely connected with the problem of exponential dichotomy of solutions. For the discussion of the bounded solutions problem from different points of view and related questions, see~\cite{Akhmerov-Kurbatov,BaskakovMS15:eng,Boichuk-Pokutnii,Chicone-Latushkin99,
Daletskii-Krein:eng,
Godunov94:ODE:eng,Hartman02:eng,Henry81:eng,
Massera-Schaffer:eng,
Pankov90:eng,Pechkurov12:eng
} and the references therein.

It is known (Theorem~\ref{t:Green}) that equation~\eqref{e:eq} has a unique bounded solution $x$ for any bounded continuous free term~$f$ if and only if the spectrum of the coefficient $A$ is disjoint from the imaginary axis. In this case, the solution can be represented in the form
\begin{equation*}
x(t)=\int_{-\infty}^{\infty}\mathcal G(A,s)f(t-s)\,ds.
\end{equation*}
The kernel $\mathcal G$ is called \emph{Green's function}.
Assuming that $A$ is triangular or a triangular form of $A$ is known, we establish an effective estimate of Green's function (Theorem~\ref{t:fin}). Estimates of Green's function are important, e.g., for numerical methods~\cite{Bulgakov80:eng,Bulgakov89a:eng,Godunov86:SMZh:eng,
Kurbatov-Kurbatova-CMAM18,Malyshev91:eng}.

\section{The definition of Green's function}\label{s:Green's function}
For $\lambda\in\mathbb C$ and $t\in\mathbb R$, we consider the functions
\begin{align*}
\exp_{t}^+(\lambda)&=\begin{cases}
e^{\lambda t}, & \text{if $t>0$},\\
0, & \text{if $t<0$},\end{cases}\\
\exp_{t}^-(\lambda)&=\begin{cases}
0, & \text{if $t>0$},\\
-e^{\lambda t}, & \text{if $t<0$},\end{cases}
\\
g_t(\lambda)&=\begin{cases}
\exp_{t}^-(\lambda), & \text{if $\Real\lambda>0$},\\
\exp_{t}^+(\lambda), & \text{if $\Real\lambda<0$}.
\end{cases}
\end{align*}
These functions are undefined for $t=0$. The function $g_t$ is also undefined for $\Real\lambda=0$. For any fixed $t\neq0$, all three functions are analytic on their domains.

Let $A$ be a complex matrix of the size $n\times n$.
We consider the differential equation
\begin{equation}\label{e:x'=Ax+f}
x'(t)=Ax(t)+f(t),\qquad t\in\mathbb R.
\end{equation}
The \emph{bounded solutions problem} is the problem of finding bounded solution $x:\,\mathbb R\to\mathbb C^n$ under the assumption that the free term $f:\,\mathbb R\to\mathbb C^n$ is a bounded function.

\begin{theorem}[{\rm\cite[Theorem 4.1, p.~81]{Daletskii-Krein:eng}}]\label{t:Green}
Let $A$ be a complex matrix of the size $n\times n$.
Equation~\eqref{e:x'=Ax+f} has a unique bounded on $\mathbb R$ solution $x$ for any bounded continuous function $f$ if and only if the spectrum $\sigma(A)$ of $A$ does not intersect the imaginary axis.
This solution admits the representation
\begin{equation*}
x(t)=\int_{-\infty}^\infty \mathcal G(A,s)f(t-s)\,ds,
\end{equation*}
where
\begin{equation*}
\mathcal G(A,t)=\frac1{2\pi i}\int_\Gamma g_t(\lambda)(\lambda\mathbf1-A)^{-1}\,d\lambda,\qquad t\neq0,
\end{equation*}
the contour $\Gamma$ encloses $\sigma(A)$, and $\mathbf1$ is the identity matrix.
\end{theorem}
The function $\mathcal G$ is called~\cite{Daletskii-Krein:eng} \emph{Green's function} of the bounded solutions problem for equation~\eqref{e:x'=Ax+f}.

\begin{corollary}\label{c:A=B}
Let $A=Q^{-1}BQ$, where $Q$ is unitary. Then
\begin{equation*}
\lVert\mathcal G(A,t)\rVert_{2\to2}=\lVert\mathcal G(B,t)\rVert_{2\to2},\qquad t\neq0,
\end{equation*}
where $\lVert\cdot\rVert_{2\to2}$ is the matrix norm induced by the norm $\lVert z\rVert_2=\sqrt{|z_1|^2+\dots+|z_n|^2}$ on $\mathbb C^n$.
\end{corollary}
\begin{proof}
It is well known that for any analytic function $f$
\begin{equation*}
f(A)=Q^{-1}f(B)Q.\qed
\end{equation*}
\renewcommand\qed{}
\end{proof}

\section{The estimate}\label{s:estimate}

\begin{proposition}\label{p:ini}
For any square matrices $A$ and $B$ of the same size with the spectrum disjoint from the imaginary axis
\begin{equation*}
\mathcal G(A,t)-\mathcal G(B,t)=\int_{-\infty}^{+\infty}\mathcal G(A,t)(A-B)\mathcal G(B,t)\,dt.
\end{equation*}
\end{proposition}
\begin{proof}
We organize the proof as a sequence of references. By~\cite[Corollary 47]{Kurbatov-Kurbatova-Oreshina},
\begin{equation*}
\mathcal G(A,t)-\mathcal G(B,t)=\frac1{2\pi i}\int_\Gamma g_t(\lambda)(\lambda\mathbf1-A)^{-1}(A-B)(\lambda\mathbf1-B)^{-1}\,d\lambda,
\end{equation*}
where the contour $\Gamma$ surrounds $\sigma(A)\cup\sigma(B)$. By~\cite[Theorem 45]{Kurbatov-Kurbatova-Oreshina},
\begin{multline*}
\frac1{2\pi i}\int_\Gamma g_t(\lambda)(\mathbf1-A)^{-1}(A-B)(\mathbf1-B)^{-1}\,d\lambda\\=
\frac1{(2\pi i)^2}\int_{\Gamma_1}\int_{\Gamma_2} g_t^{[1]}(\lambda,\mu)(\lambda\mathbf1-A)^{-1}(A-B)(\mu\mathbf1-B)^{-1}\,d\mu\,d\lambda,
\end{multline*}
where $g_t^{[1]}$ is the divided difference of the function $g_t$, $\Gamma_1$ surrounds $\sigma(A)$, and $\Gamma_2$ surrounds $\sigma(B)$. We recall that the \emph{divided difference}~\cite{Gelfond:eng,Jordan} of a function $f$ is the function
\begin{equation}\label{e:f[1]}
f^{[1]}(\lambda,\mu)= \begin{cases}
\frac{f(\lambda)-f(\mu)}{\lambda-\mu}, & \text{if $\lambda\neq\mu$},\\
f'(\lambda), & \text{if $\lambda=\mu$}.
 \end{cases}
\end{equation}
For more about $g_t^{[1]}$, see, e.g.,~\cite{Kurbatov-Kurbatova-CMAM18}.
By~\cite[Theorem 23]{Kurbatov-Kurbatova-triangular1},
\begin{multline*}
\frac1{(2\pi i)^2}\int_{\Gamma_1}\int_{\Gamma_2} g_t^{[1]}(\lambda,\mu)(\lambda\mathbf1-A)^{-1}(A-B)(\mu\mathbf1-B)^{-1}\,d\mu\,d\lambda\\
=\int_{-\infty}^{\infty}\mathcal G(A,s)(A-B)\mathcal G(B,t-s)\,ds.\qed
\end{multline*}
\renewcommand\qed{}
\end{proof}

The main result of this paper is the following theorem.

\begin{theorem}\label{t:fin}
Let a complex triangular matrix $B$ have the size $n\times n$ and be represented as the sum $B=D+N$ of a diagonal matrix $D$ and a strictly triangular matrix $N$. Let $\gamma^->0$ and $\gamma^+>0$ be chosen so that the strip $\{z\in\mathbb C:\,-\gamma^+<\Real z<\gamma^-\,\}$ is disjoint from the spectrum $\sigma(B)$ of the matrix $B$.
Let the norm on the space of matrices be induced by a norm on $\mathbb C^n$ and possesses the property $\lVert D\rVert\le\max_i|d_{ii}|$ for any diagonal matrix $D$.
Then
\begin{equation}\label{e:est}
\lVert \mathcal G(B,t)\rVert
\le\bigl(\heaviside(t)e^{-\gamma^-t}+\heaviside(-t)e^{\gamma^+t}\bigr)\sum_{k=0}^{n-1}\lVert N\rVert^{k}
\frac{|t|^{k}
\sqrt{\gamma|t|}\,e^{\gamma|t|/2} K_{-k-\frac{1}{2}}\left(\frac{1}{2}
\gamma|t|\right)}{\sqrt{\pi}k!},
\end{equation}
where $K$ is the modified Bessel function of the second kind~\cite{Watson95:eng}, $\heaviside$ is the Heaviside function
\begin{equation*}
\heaviside(t)=
\begin{cases}
1& \text{ if $t>0$}, \\
0& \text{ if $t<0$}
\end{cases}
\end{equation*}
and
\begin{equation*}
\gamma=\gamma^-+\gamma^+.
\end{equation*}
\end{theorem}

\begin{remark}\label{r:2->2}
With the help of Corollary~\ref{c:A=B}, in the case of the $\lVert\cdot\rVert_{2\to2}$ norm, Theorem~\ref{t:fin} can be applied to any matrix $A$ provided its triangular representation $A=Q^{-1}BQ$ is known.
\end{remark}

\begin{remark}\label{r:BesselK}
It can be shown that the function $t\mapsto\dfrac{t^{k}\sqrt{\gamma t}\,e^{\gamma t/2} K_{-k-\frac{1}{2}}\left(\frac{1}{2}
\gamma t\right)}{\sqrt{\pi}k!}$ is a polynomial of degree $k$ with the leading term $\dfrac{t^{k}}{k!}$. For example, if $B$ is normal, then $N=0$ and, thus,
\begin{equation*}
\sum_{k=0}^{n-1}\lVert N\rVert^{k}\frac{|t|^{k}
\sqrt{\gamma|t|}\,e^{\gamma|t|/2} K_{-k-\frac{1}{2}}\left(\frac{1}{2}
\gamma|t|\right)}{\sqrt{\pi}k!}=\lVert N\rVert^{0}
\frac{|t|^{0}
\sqrt{\gamma|t|}\,e^{\gamma|t|/2} K_{-\frac{1}{2}}\left(\frac{1}{2}
\gamma|t|\right)}{\sqrt{\pi}0!}=1.
\end{equation*}
\end{remark}

\begin{remark}\label{r:Vin-Gro}
Another similar estimate of Green's function is proved in~\cite{Kurbatov-Kurbatova-QTDS18}:
\begin{align*}
\Vert\mathcal G(A,t)\Vert&\le e^{-\gamma^-t}\sum_{j=0}^{m-1}
\sum_{i=0}^j\binom{l+i-1}{l-1}\frac{t^{j-i}}{(j-i)!}
\frac{(2\Vert A\Vert)^{l+j}}{\gamma^{l+i}},& t&>0,\\
\Vert\mathcal G(A,t)\Vert&\le e^{\gamma^+t}\sum_{j=0}^{l-1}
\sum_{i=0}^j\binom{m+i-1}{m-1}\frac{t^{j-i}}{(j-i)!}
\frac{(2\Vert A\Vert)^{m+j}}{\gamma^{m+i}},& t&<0.
\end{align*}
Here $m$ (respectively, $l$) is the number of eigenvalues of $A$ counted according to their algebraic multiplicities that lie in the open left (right) half-plane. In this estimate the maximal powers of $t$ are $m-1$ and $l-1$ whereas the highest power of $t$ in~\eqref{e:est} is $n-1$. On the other hand, estimate from~\cite{Kurbatov-Kurbatova-QTDS18} contains the factors $\lVert A\rVert^{k}$ instead of $\lVert N\rVert^{k}$ in~\eqref{e:est}, which may be essentially smaller; for example, $N=0$ if $A$ is normal.
\end{remark}

\begin{proof}[Proof of Theorem~\ref{t:fin}]
From Proposition~\ref{p:ini} it follows that
\begin{equation}\label{e:G(D+N):1}
\mathcal G(B,t)=\mathcal G(D+N,t)=\mathcal G(D,t)+\int_{-\infty}^{\infty}\mathcal G(D+N,s)N\mathcal G(D,t-s)\,ds.
\end{equation}
We substitute this representation into itself:
\begin{align*}
\mathcal G(B,t)&=\mathcal G(D,t)+\int_{-\infty}^{\infty}
\mathcal G(D,s_1)N\mathcal G(D,t-s_1)\,ds_1\\
&+
\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\mathcal G(D+N,s_2)N\mathcal G(D,s_1-s_2)
N\mathcal G(D,t-s_1)\,ds_2\,ds_1.
\end{align*}
Then we substitute~\eqref{e:G(D+N):1} again into the obtained formula several times:
\begin{equation}\label{e:G(B,t)=}
\begin{split}
\mathcal G(B,t)&=\mathcal G(D,t)+\int_{-\infty}^{\infty}
\mathcal G(D,s_1)N\mathcal G(D,t-s_1)\,ds_1\\
&+
\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\mathcal G(D,s_2)N\mathcal G(D,s_1-s_2)
N\mathcal G(D,t-s_1)\,ds_2\,ds_1\\
&+
\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\mathcal G(D,s_3)N\mathcal G(D,s_2-s_3)\\
&\times N\mathcal G(D,s_1-s_2)
N\mathcal G(D,t-s_1)\,ds_3\,ds_2\,ds_1+\dots\\
&+
\int_{-\infty}^{\infty}\dots\int_{-\infty}^{\infty}\mathcal G(D,s_{n-1})N\mathcal G(D,s_{n-2}-s_{n-1})\dots\\
&\times N\mathcal G(D,s_1-s_2)
N\mathcal G(D,t-s_1)\,ds_{n-1}\dots\,ds_1.
 \end{split}
\end{equation}
The subsequent terms are zero because of the following reason.
Since the matrix $D$ is diagonal, the matrix function $\mathcal G(D,\cdot)$ is also diagonal. The matrix $N$ is strictly triangular. Therefore the diagonal of $N\mathcal G(\cdot\cdot)N\mathcal G(\cdot)$ closest to the main one is zero, and so on.

From representation~\eqref{e:G(B,t)=} it follows that
\begin{align*}
\lVert \mathcal G(B&,t)\rVert\le\lVert\mathcal G(D,t)\rVert+
\lVert N\rVert\int_{-\infty}^{\infty}
\lVert \mathcal G(D,s_1)\rVert\lVert \mathcal G(D,t-s_1)\rVert\,ds_1\\
&+
\lVert N\rVert^2
\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\lVert \mathcal G(D,s_2)\rVert
\lVert \mathcal G(D,s_1-s_2)\rVert
\lVert \mathcal G(D,t-s_1)\rVert\,ds_2\,ds_1+\dots\\
&+
\lVert N\rVert^{n-1}\int_{-\infty}^{\infty}\dots\int_{-\infty}^{\infty}\lVert \mathcal G(D,s_{n-1})\rVert
\lVert \mathcal G(D,s_{n-2}-s_{n-1})\rVert\dots\\
&\times \lVert \mathcal G(D,s_1-s_2)\rVert
\lVert \mathcal G(D,t-s_1)\rVert\,ds_{n-1}\dots\,ds_1.
\end{align*}

We set
\begin{equation*}
h(t)=
\begin{cases}
e^{-\gamma^-t} & \text{ for $t\ge0$ }, \\
e^{\gamma^+t} & \text{ for $t\le0$ }.
\end{cases}
\end{equation*}
Since the matrix $D$ is diagonal,
\begin{equation*}
\lVert\mathcal G(D,t)\rVert\le h(t).
\end{equation*}
Therefore,
\begin{align*}
\lVert \mathcal G(D&+N,t)\rVert\le h(t)+
\lVert N\rVert\int_{-\infty}^{\infty}h(s_1)h(t-s_1)\,ds_1\\
&+
\lVert N\rVert^2
\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}h(s_2)h(s_1-s_2)h(t-s_1)\,ds_2\,ds_1+\dots\\
&+
\lVert N\rVert^{n-1}\int_{-\infty}^{\infty}\dots\int_{-\infty}^{\infty}
h(s_{n-1})h(s_{n-2}-s_{n-1})\dots h(t-s_1)\,ds_{n-1}\dots\,ds_1.
\end{align*}
We denote the $k$-fold convolution of $h$ with itself by $h^{*k}$:
\begin{equation*}
h^{*k}(t)=\int_{-\infty}^{\infty}\dots\int_{-\infty}^{\infty}
h(s_{k-1})h(s_{k-2}-s_{k-1})\dots h(t-s_1)\,ds_{k-1}\dots\,ds_1.
\end{equation*}
With this notation, the previous estimate takes the from
\begin{equation*}
\lVert\mathcal G(B,t)\rVert\le\sum_{k=0}^{n-1}\lVert N\rVert^{k}h^{*(k+1)}(t).
\end{equation*}

Next we calculate $h^{*k}(t)$. We do some calculations with the help of `Mathematica'~\cite{Wolfram}.

Clearly, the Fourier transform $\hat h$ of $h$ is
\begin{equation*}
\hat h(\omega)=\frac1{\gamma^-+i\omega}+\frac1{\gamma^+-i\omega}.
\end{equation*}
Therefore the Fourier transform $\widehat{h^{*k}}$ of $h^{*k}$ is
\begin{equation*}
\widehat{h^{*k}}(\omega)=h^k(\omega)=\Bigl(\frac1{\gamma^-+i\omega}+\frac1{\gamma^+-i\omega}\Bigr)^k
=\sum_{j=0}^k\binom{k}{j}\frac1{(\gamma^+-i\omega)^j}\frac1{(\gamma^-+i\omega)^{k-j}}.
\end{equation*}

Further we have
\begin{align*}
\frac1{(\gamma^+-i\omega)^j}\frac1{(\gamma^-+i\omega)^{k-j}}&=
\sum_{q=0}^{k-j-1}\binom{j+q-1}{q}\frac1{\gamma^{j+q}(\gamma^-+i\omega)^{k-j-q}}\\
&+\sum_{q=0}^{j-1}\binom{k-j+q-1}{q}\frac1{\gamma^{k-j+q}(\gamma^+-i\omega)^{j-q}}.
\end{align*}
Consequently,
\begin{align*}
\widehat{h^{*k}}(\omega)
&=\sum_{j=0}^k\binom{k}{j}\frac1{(\gamma^+-i\omega)^j}\frac1{(\gamma^-+i\omega)^{k-j}}\\
&=\sum_{j=0}^k\binom{k}{j}
\sum_{q=0}^{k-j-1}\binom{j+q-1}{q}\frac1{\gamma^{j+q}(\gamma^-+i\omega)^{k-j-q}}\\
&+\sum_{j=0}^k\binom{k}{j}\sum_{q=0}^{j-1}\binom{k-j+q-1}{q}\frac1{\gamma^{k-j+q}(\gamma^+-i\omega)^{j-q}}\\
\end{align*}

It is well known that the inverse Fourier transform takes the functions
\begin{equation*}
\omega\mapsto\frac1{(\gamma^-+i\omega)^{k-j-q}},\qquad
\omega\mapsto\frac1{(\gamma^+-i\omega)^{j-q}},
\end{equation*}
respectively, to the functions
\begin{equation*}
t\mapsto\heaviside(t)\,\frac{t^{k-j-q-1}}{(k-j-q-1)!}e^{-\gamma^-t},\qquad
t\mapsto-\heaviside(-t)\,\frac{(-1)^{j-q}t^{j-q-1}}{(j-q-1)!}e^{\gamma^+t}.
\end{equation*}
Hence
\begin{align*}
h^{*k}(t)&=e^{-\gamma^-t}\,\heaviside(t)\sum_{j=0}^k\binom{k}{j}
\sum_{q=0}^{k-j-1}\binom{j+q-1}{q}\frac1{\gamma^{j+q}}
\frac{t^{k-j-q-1}}{(k-j-q-1)!}\\
&-e^{\gamma^+t}\,\heaviside(-t)\sum_{j=0}^k\binom{k}{j}
\sum_{q=0}^{j-1}\binom{k-j+q-1}{q}\frac1{\gamma^{k-j+q}}
\frac{(-1)^{j-q}t^{j-q-1}}{(j-q-1)!}.
\end{align*}
We change the order of summation:
\begin{align*}
h^{*k}(t)
&=e^{-\gamma^-t}\,\heaviside(t)\sum_{p=0}^{k-1}\sum_{j=0}^{p}
\binom{k}{j}\binom{p-1}{p-j}\frac1{\gamma^{p}}
\frac{t^{k-p-1}}{(k-p-1)!}\\
&-e^{\gamma^+t}\,\heaviside(-t)\sum_{p=1}^k\sum_{q=0}^{k-p}
\binom{k}{p+q}\binom{k-p-1}{q}\frac1{\gamma^{k-p}}
\frac{(-1)^{p}t^{p-1}}{(p-1)!}.
\end{align*}
First, we calculate the internal sums (here we essentially use `Mathematica'):
\begin{align*}
h^{*k}(t)
&=e^{-\gamma^-t}\,\heaviside(t)\sum_{p=0}^{k-1}
\frac{t^{k-p-1}(k+p-1)!}{\gamma^{p}\,(k-1)!p!(k-p-1)!}\\
&-e^{\gamma^+t}\,\heaviside(-t)\sum_{p=1}^k
\frac{(-1)^{k-p} t^{k-p-1}(k+p-1)!
}{\gamma^{p}\,(k-1)!p!(k-p-1)!}.
\end{align*}
Then, by calculating the final sums (here we again essentially use `Mathematica'), we arrive at
\begin{equation}\label{e:est of h}
\begin{split}
h^{*k}(t)
&=e^{-\gamma^-t}\,\heaviside(t)\frac{t^{k-1}
\sqrt{\gamma t}\,e^{\frac{1}{2}\gamma t} K_{\frac{1}{2}-k}\left(\frac{1}{2}
\gamma t\right)}{\sqrt{\pi }(k-1)!}\\
&+e^{\gamma^+t}\,\heaviside(-t)(-1)^{k-1}\frac{t^{k-1}\sqrt{-\gamma t}\,e^{-\frac{1}{2}\gamma t} K_{\frac{1}{2}-k}\left(-\frac{1}{2}\gamma
   t\right)}{\sqrt{\pi }(k-1)!}\\
&=h(t)\frac{|t|^{k-1}
\sqrt{\gamma|t|}\,e^{\gamma|t|/2} K_{\frac{1}{2}-k}\left(\frac{1}{2}
\gamma|t|\right)}{\sqrt{\pi}(k-1)!}.\qed
\end{split}
\end{equation}
\renewcommand\qed{}
\end{proof}

\begin{remark}\label{r:limit}
When $\gamma^+\to+\infty$, we have $e^{\gamma^+t}\to0$ for $t<0$ and (which is calculated in `Mathematica')
\begin{equation*}
\lim_{\gamma^+\to+\infty}\frac{|t|^{k}
\sqrt{\gamma|t|}\,e^{\gamma|t|/2} K_{-k-\frac{1}{2}}\left(\frac{1}{2}
\gamma|t|\right)}{\sqrt{\pi}k!}=1.
\end{equation*}
Therefore estimate~\eqref{e:est} turns into the estimate from~\cite{Van_Loan77}:
\begin{equation*}
\lVert\mathcal G(B,t)\rVert\le\heaviside(t) e^{-\gamma^-t}\sum_{k=0}^{n-1}\frac{\lVert N\rVert^{k}t^{k}}{k!}.
\end{equation*}
\end{remark}

Sometimes $\lVert N^k\rVert$ may decrease faster than $\lVert N\rVert^k$. In such a case it may be more convenient to use the following variant of Theorem~\ref{t:fin}.



\begin{corollary}\label{c:5}
Let the norm on the space of matrices be induced by the norm $\lVert z\rVert_1=|z_1|+\dots+|z_n|$ or the norm $\lVert z\rVert_\infty=\max_i|z_i|$ on $\mathbb C^n$. Then under the assumptions of Theorem~\ref{t:fin}
\begin{equation*}
|\mathcal G(B,t)|
\le\bigl(\heaviside(t)e^{-\gamma^-t}+\heaviside(-t)e^{\gamma^+t}\bigr)\sum_{k=0}^{n-1} |N|^{k}
\frac{|t|^{k}
\sqrt{\gamma|t|}\,e^{\gamma|t|/2} K_{-k-\frac{1}{2}}\left(\frac{1}{2}
\gamma|t|\right)}{\sqrt{\pi}k!},
\end{equation*}
where $|A|$ is the matrix consisting of absolute values $|a_{ij}|$ of the entries $a_{ij}$ of a matrix $A$, and the inequality $A\le B$ for matrices means the corresponding inequalities for their entries.
\end{corollary}

\begin{remark}\label{r:Loan}
The idea of using the power $|N|^k$ was used in~\cite[Theorem 11.2.2]{Golub-Van_Loan96:eng} for the estimating of the general function of a matrix.
\end{remark}

\begin{proof}
Both the norms in question possess the evident properties: $|A+B|\le|A|+|B|$, $|A\cdot B|\le|A|\cdot|B|$, $\lVert A\rVert=\lVert |A|\rVert$, and $\lVert |A|\rVert\le\lVert|B|\rVert$ if $|A|\le|B|$. Hence, from~\eqref{e:G(B,t)=} we have
\begin{align*}
|\mathcal G(B,t)|&\le|\mathcal G(D,t)|+\int_{-\infty}^{\infty}
|\mathcal G(D,s_1)N\mathcal G(D,t-s_1)|\,ds_1\\
&+
\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}|\mathcal G(D,s_2)N\mathcal G(D,s_1-s_2)
N\mathcal G(D,t-s_1)|\,ds_2\,ds_1\\
&+
\int_{-\infty}^{\infty}\dots\int_{-\infty}^{\infty}|\mathcal G(D,s_{n-1})N\mathcal G(D,s_{n-2}-s_{n-1})\dots\\
&\times N\mathcal G(D,s_1-s_2)
N\mathcal G(D,t-s_1)|\,ds_{n-1}\dots\,ds_1.
\end{align*}
Therefore,
\begin{align*}
|\mathcal G(B,t)|&\le h(t)\mathbf1+\int_{-\infty}^{\infty}
(h(s_1)\mathbf1)|N|(h(t-s_1)\mathbf1)\,ds_1\\
&+
\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}(h(s_2)\mathbf1)|N|(h(s_1-s_2)\mathbf1)
|N|(h(t-s_1)\mathbf1)\,ds_2\,ds_1\\
&+
\int_{-\infty}^{\infty}\dots\int_{-\infty}^{\infty} (h(s_{n-1})\mathbf1)|N|(h(s_{n-2}-s_{n-1})\mathbf1)\dots\\
&\times |N|(h(s_1-s_2)\mathbf1)
|N|(h(t-s_1)\mathbf1)\,ds_{n-1}\dots\,ds_1.
\end{align*}
Or
\begin{align*}
|\mathcal G(B,t)|&\le h(t)\mathbf1+\int_{-\infty}^{\infty}h(s_1)h(t-s_1)
|N|\,ds_1\\
&+
\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}h(s_2)h(s_1-s_2)h(t-s_1)
|N|^2\,ds_2\,ds_1\\
&+
\int_{-\infty}^{\infty}\dots\int_{-\infty}^{\infty}h(s_{n-1})(h(s_{n-2}-s_{n-1})\dots h(s_1-s_2)|N|^{n-1}\,ds_{n-1}\dots\,ds_1\\
&=\sum_{k=0}^{n-1}|N|^{k} h^{*(k+1)}(t).
\end{align*}
Now the proof follows from estimate~\eqref{e:est of h}.
\end{proof}

\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}

\end{document}